OpenStudy (anonymous):

Which equation is a decomposition reaction? Li2CO3 mc016-1.jpg Li2O + CO2 Zn + HCl mc016-2.jpg ZnCl2 + H2 Na2O + CO2 mc016-3.jpg Na2CO3 C6H12O6 + O2 mc016-4.jpg CO2 + H2O @sweetburger

OpenStudy (anonymous):

is it the 2nd one

OpenStudy (anonymous):

Li2CO3 = Li2O + CO2 Zn + HCl = ZnCl2 + H2 Na2O + CO2 = Na2CO3 C6H12O6 + O2= CO2 + H2O

OpenStudy (sweetburger):

never heard of the compound mc016-1.jpg must be pretty fancy

OpenStudy (anonymous):

i fixed it

OpenStudy (sweetburger):

lol

OpenStudy (sweetburger):

remember its the breaking apart of a compound or molecule so which reaction does this?

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

C6H12O6 + O2= CO2 + H2O

OpenStudy (sweetburger):

no thats a combustion reaction as you see its a hydrocarbon reacting with O2 to produce H2O and CO2

OpenStudy (sweetburger):

Its going to be 1 reactant breaking into multiple products

OpenStudy (anonymous):

so Na2O + CO2 = Na2CO3?

OpenStudy (anonymous):

then it must be Li2CO3 = Li2O + CO2

OpenStudy (anonymous):

it has multiple products

OpenStudy (sweetburger):

yes you are correct

OpenStudy (anonymous):

Li2CO3 = Li2O + CO2 is the answer

OpenStudy (sweetburger):

yes